WACC AND OPTIMAL CAPITAL BUDGET Adamson Corporation is considering four average-risk projects with the following costs and rates of return: Project Cost Expected Rate of Return 1 $2,000 16.00% 2 3,000 15.00 3 5,000 13.75 4 2,000 12.50 The company estimates that it can issue debt at a rate of rd = 10%, and its tax rate is 30%. It can issue preferred stock that pays a constant dividend of $5 per year at $54 per share. Also, its common stock currently sells for $32 per share; the next expected dividend, D1, is $3.75; and the dividend is expected to grow at a constant rate of 5% per year. The target capital structure consists of 75% common stock, 15% debt, and 10% preferred stock.

What is the cost of each of the capital components? Round your answers to two decimal places. Do not round your intermediate calculations.

Cost of debt _____%

Cost of preferred stock _____%

Cost of retained earnings _____%

What is Adamson's WACC? Round your answer to two decimal places. Do not round your intermediate calculations.

______%
Answer #1

1.
=10%*(1-30%)=7.00%

2.
=5/54
=9.2593%

3.
=3.75/32+5%
=16.7188%

4.
=75%*16.7188%+15%*7.00%+10%*9.2593%
=14.5150%